# Bounding User Contributions for User-Level Differentially Private Mean Estimation

### v3

@@ Abstract
We revisit the problem of releasing the sample mean of bounded samples in a dataset, privately, under user-level $\varepsilon$-differential privacy (DP). We aim to derive the optimal method of preprocessing data samples, within a canonical class of processing strategies, in terms of the error in estimation. Typical error analyses of such bounding (or clipping) strategies in the literature assume that the data samples are independent and identically distributed (i.i.d.), and sometimes also that all users contribute the same number of samples (data homogeneity)—assumptions that do not accurately model real-world data distributions. Our main result in this work is a precise characterization of the preprocessing strategy that gives rise to the smallest worst-case error over all datasets – a distribution-independent error metric – while allowing for data heterogeneity. We also show via experimental studies that even for i.i.d. real-valued samples, our clipping strategy performs much better, in terms of average-case error, than the widely used bounding strategy of Amin et al. (2019).
@@ I Introduction
In this article, we concern ourselves with the fundamental problem of processing bounded, potentially vector-valued samples in a dataset, for the release of a private estimate of the sample mean. In particular, we work within the framework of “user-level” differential privacy [1, 2], which is a generalization of the now widely adopted framework of differential privacy (DP) [3, 4] for the design and analysis of privacy-preserving algorithms. Loosely speaking, user-level DP guarantees the privacy of a “user”, who could contribute more than one sample, by ensuring the statistical indistinguishability of outputs of the algorithm to changes in the user’s samples. User-level DP has practical relevance for inference tasks on most real-world datasets, such as traffic datasets, datasets of user expenditures, and time series data, where different users contribute potentially different numbers of samples (data heterogeneity) [5, 6]. Moreover, user-level DP algorithms are increasingly becoming popular subroutines for integration into federated learning (FL) frameworks – see, e.g., [7, Sec. 4] for more details.
@@
There are two key requirements of such user-level DP mechanisms for mean estimation, for real-world applications. Firstly, the mechanisms must be designed to work with heterogeneous data. Secondly, one would like reliable reconstruction of the true sample mean, even when the data samples are non-i.i.d. (independent and identically distributed). Our focus is hence to characterize an error metric, which is independent of the underlying data distribution and can be explicitly computed and optimized, for heterogeneous data.
@@
In this article, we confine our attention to (pure) $\varepsilon$-DP algorithms for mean estimation. A key subroutine in most user-level $\varepsilon$-DP mechanisms [1, 8, 9, 10] is the preprocessing of the data samples for the release of an estimate of the sample mean, which requires the addition of less noise for privacy (measured via the sensitivity of the estimate), as against releasing a noised version of the true sample mean. Such a preprocessing procedure (also called a strategy for “bounding” or “clipping” strategy [10]) either drops certain samples contributed by selected users, or projects the samples to a “high-probability interval” that is a strict subset of the interval in which the sample values are known to lie. While it is usually easy to establish that the mechanisms designed using the clipped estimators are differentially private, an analysis of their “utility”, or the error in estimation of the true statistic, often relies on distributional assumptions about the dataset.
@@
In this work, following [11, 12, 13], we define and explicitly compute the worst-case error, over all datasets, of general preprocessing (or bounding) strategies. The worst-case error metric is natural in settings with arbitrarily correlated data, where each user potentially ascribes his/her error tolerance to the worst dataset that the statistic is computed on. Furthermore, this error metric is clearly distribution independent and is computable under data heterogeneity too. We then explicitly identify the bounding strategy that results in the smallest worst-case error; our approach is hence an extension of those in [10] to the setting of the worst-case error, while allowing for arbitrary bounding strategies. Interestingly, we also observe from experimental studies that for scalar samples, our clipping strategy also gives rise to much smaller errors on average compared to the strategy in [10], for selected dataset sizes, when the samples are drawn i.i.d. according to common distributions.
@@ II-B Problem Formulation
Let $L$ be the number of users present in the dataset. For every user $\ell\in[L]$, let the number of records contributed by the user be denoted by $m_{\ell}$, and set $m^{\star}:=\max_{\ell\in[L]}m_{\ell}$ and $m_{\star}:=\min_{\ell\in[L]}m_{\ell}$. We assume that $L$ and the collection $\{m_{\ell}:\ell\in[L]\}$ are known. Now, for a given user $\ell\in[L]$, let $\left\{\mathbf{x}_{j}^{(\ell)}:j\in[m_{\ell}]\right\}$ denote the collection of (potentially arbitrary) bounded samples contributed by the user, where each $\mathbf{x}_{j}^{(\ell)}\in\mathbb{R}^{d}$, for some dimension $d\geq 1$. We assume, as is common for most applications, that $\left\lVert\mathbf{x}_{j}^{(\ell)}\right\rVert_{1}\leq U$ [14, 11], for all $\ell\in[L]$, $j\in[m_{\ell}]$, where $U>0$ is known. Call the dataset as $\mathcal{D}=\left\{\left(\ell,\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$.
We are interested in releasing the sample mean $f=f(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{x}_{j}^{(\ell)}.$
@@ II-C Differential Privacy
Consider datasets $\mathcal{D}_{1}=\left\{\left(u_{\ell},\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$ and $\mathcal{D}_{2}=\left\{\left(u_{\ell},\overline{\mathbf{x}}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$ consisting of the same users, with each user contributing the same number, $m_{\ell}$, of data values. Let $\mathsf{D}$ denote a universal set of such datasets. We say that $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are “user-level neighbours” if there exists $\ell_{0}\in[L]$ such that $\left(\mathbf{x}^{(\ell_{0})}_{1},\ldots,\mathbf{x}^{(\ell_{0})}_{m_{\ell_{0}}}\right)\neq\left(\overline{\mathbf{x}}^{(\ell_{0})}_{1},\ldots,\overline{\mathbf{x}}^{(\ell_{0})}_{m_{\ell_{0}}}\right)$, with $\left(\mathbf{x}^{(\ell)}_{1},\ldots,\mathbf{x}^{(\ell)}_{m_{\ell}}\right)=\left(\overline{\mathbf{x}}^{(\ell)}_{1},\ldots,\overline{\mathbf{x}}^{(\ell)}_{m_{\ell}}\right)$, for all $\ell\neq\ell_{0}$.
@@
We use the terms “sensitivity” and “user-level sensitivity” interchangeably. The next result is well-known [15, Prop. 1].
@@ II-D The Worst-Case Error Metric
All through, in this paper, we shall work with user-level $\varepsilon$-DP mechanisms that add a suitable amount of Laplace noise that is tailored to the sensitivity of the function used as an estimator of the sample mean $f$. Consider a mechanism $M$ for the user-level $\varepsilon$-DP release of the statistic $f$. The canonical structure of $M$ (see [16, Footnote 1], [17]) is: $M(\mathcal{D})=\overline{f}(\mathcal{D})+{\mathbf{Z}},$ for some estimator $\overline{f}$ of $f$, with user-level sensitivity $\Delta_{\overline{f}}$, with $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$, with $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}\left(\Delta_{\overline{f}}/\varepsilon\right)$, $i\in[d]$.
@@
$E_{M}=E_{\overline{f}}:=\max_{\mathcal{D}\in\mathsf{D}}\big{\lVert}f(\mathcal{D})-\overline{f}(\mathcal{D})\big{\rVert}_{1}+\mathbb{E}[\lVert{\mathbf{Z}}\rVert_{1}].$ (1)
Clearly, the expression above is an upper bound on the $\ell_{1}$-error $E^{(1)}_{M}:=\max_{\mathcal{D}\in\mathsf{D}}\mathbb{E}[|f(\mathcal{D})-f(\mathcal{D})|]$, via the triangle inequality. The distribution-independent expression in (1) conveniently captures the errors due to bias (the first term) and due to noise addition for privacy (the second term); a similar such error measure that separates the bias and noise errors was employed in [10].
III Worst-Case Errors of Bounding Strategies
@@ III-A On Clipping Strategies
We work with estimators $\overline{f}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ of $f$ obtained by bounding user contributions as follows: for each $\ell\in[L]$ and $j\in[m_{\ell}]$, we let $\overline{\mathbf{x}}_{j}^{(\ell)}:={\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}}\left(\mathbf{x}_{j}^{(\ell)}\right)$, for reals $0\leq a_{j}^{(\ell)}\leq b_{j}^{(\ell)}\leq U$. In words, $\overline{\mathbf{x}}_{j}^{(\ell)}$ is a projection of $\mathbf{x}_{j}^{(\ell)}$ onto the set $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$, which, intuitively, reduces the range of values that $\mathbf{x}_{j}^{(\ell)}$ can take, and hence its sensitivity too. Figure 1 shows a pictorial depiction of the annulus $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ when $d=2$ and examples of projections onto the annulus. We then set
@@
Note that the class $\mathsf{B}$ of estimators $\overline{f}$ as above captures those estimators obtained by dropping selected samples $\mathbf{x}_{j}^{(\ell)}$ (by setting $a_{j}^{(\ell)}=b_{j}^{(\ell)}=0$ for those samples) and those obtained by projecting samples onto an $\ell_{1}$-bounded subset of $\Delta_{U}$, in addition to strategies that perform a combination of dropping and projection. This class of estimators hence includes several common estimators of the sample mean used in works such as [1, 12, 13].
@@ III-B On Worst-Case Errors of Clipping Strategies
Consider the quantity $E_{\overline{f}}$, for $\overline{f}\in\mathsf{B}$. The following proposition then holds:
@@
$E_{\overline{f}}$ $=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
The proof of Proposition III.1 proceeds with help from a few lemmas. For $E_{\overline{f}}$ as in (1), we define $\beta(\overline{f})$ to be the bias, i.e., $\beta(\overline{f}):=\max_{\mathcal{D}\in\mathsf{D}}\big{\lVert}f(\mathcal{D})-\overline{f}(\mathcal{D})\big{\rVert}$ and $\eta(\overline{f})$ to be the error due to noise addition, i.e., $\eta(\overline{f}):=\mathbb{E}[\lVert\mathbf{\mathbf{Z}}\rVert_{1}]$. First, we aim to characterize $\beta(\overline{f})$. To this end, we first state a necessary condition for a vector $\mathbf{y}$ to be an $\ell_{1}$-projection of a vector $\mathbf{a}\in\Delta_{U}$ onto $\Delta_{\alpha}$, for $\alpha\leq U$.
@@
In the appendix, we explicitly characterize the vectors $\mathbf{y}\in\Delta_{\alpha}$ that are the $\ell_{1}$-projections of $\mathbf{a}\in\Delta_{U}\setminus\Delta_{\alpha}$, for $\alpha\leq U$; indeed, we have that the condition stated in Lemma III.1 is both necessary and sufficient. The next lemma exactly characterizes $\beta(\overline{f})$.
@@
We have that $\beta(\overline{f})=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)$.
@@
We have that $\eta(\overline{f})=\frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
@@
Via arguments entirely analogous to that in the proof of Lemma III.2, the user-level sensitivity of $\overline{f}$ is
@@
since in the worst-case, all samples of a user $\ell$ are changed each from $(a_{j}^{(\ell)},0,\ldots,0)$ to $(b_{j}^{(\ell)},0,\ldots,0)$. Using $\mathbb{E}[\lVert\mathbf{Z}\rVert_{1}]=d\mathbb{E}[|Z_{1}|]=d\Delta_{\overline{f}}/\varepsilon$ gives us the required result. ∎
@@
We have that $f^{\star}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizes $E_{\overline{f}}$, where $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{Um_{\ell}+T_{\varepsilon}}{2m_{\ell}},U\right\},$ for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Some remarks are in order. First, note that the optimal bounding strategy $f^{\star}$ clips sample values based only on the number of contributions of each user $\ell\in[L]$. Furthermore, note that the interval of projection is determined by $T_{\varepsilon}$, which is quite similar in structure to the optimal clipping threshold $T$ in [10] for item-level DP, which is the (privately estimated) $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample value.
@@
There exists an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$, which obeys $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.

### v4

@@ Abstract
We revisit the problem of releasing the sample mean of bounded samples in a dataset, privately, under user-level $\varepsilon$-differential privacy (DP). We aim to derive the optimal method of preprocessing data samples, within a canonical class of processing strategies, in terms of the estimation error. Typical error analyses of such bounding (or clipping) strategies in the literature assume that the data samples are independent and identically distributed (i.i.d.), and sometimes also that all users contribute the same number of samples (data homogeneity)—assumptions that do not accurately model real-world data distributions. Our main result in this work is a precise characterization of the preprocessing strategy that gives rise to the smallest worst-case error over all datasets – a distribution-independent error metric – while allowing for data heterogeneity. We also show via experimental studies that even for i.i.d. real-valued samples, our clipping strategy performs much better, in terms of average-case error, than the widely used bounding strategy of Amin et al. (2019).
@@ I Introduction
In this article, we work on the fundamental problem of processing bounded, potentially vector-valued samples in a dataset, for the release of a private estimate of the sample mean. In particular, we work within the framework of “user-level” differential privacy [userlevel, hetero], which is a generalization of the now widely adopted framework of differential privacy (DP) [dworkroth, vadhan2017] for the design and analysis of privacy-preserving algorithms. Loosely, user-level DP guarantees the privacy of a “user”, who could contribute more than one sample, by ensuring the statistical indistinguishability of outputs of the algorithm to changes in the user’s samples. User-level DP has practical relevance for inference tasks on real-world datasets, such as traffic datasets, datasets of user expenditures, and time series data, where different users contribute potentially different numbers of samples (data heterogeneity) [usereg3, usereg4], and more recently, in federated learning (FL) applications (see, e.g., [kairouz-survey, Sec. 4]).
@@
There are two key requirements of such user-level DP mechanisms for mean estimation. Firstly, the mechanisms must be designed to work with heterogeneous data. Secondly, one would like reliable reconstruction of the true sample mean, even when the data samples are non-i.i.d. (independent and identically distributed). Our focus is hence to characterize an error metric, which is independent of the underlying data distribution and can be explicitly computed and optimized, for heterogeneous data.
@@
In this article, we confine our attention to user-level (pure) $\varepsilon$-DP algorithms for mean estimation. A key subroutine in such mechanisms [userlevel, hetero-user-level, hist-user-level, amin] is the preprocessing of the data samples for the release of an estimate of the sample mean, which requires the addition of less noise for privacy, as against releasing a noised version of the true sample mean. Such a preprocessing procedure (or “bounding” or “clipping” strategy [amin]) either drops selected samples, or projects the samples to a “high-probability subset”. While it is usually easy to establish that such mechanisms are differentially private, an analysis of their “utility”, or the error in estimation of the true statistic, often relies on distributional assumptions about the dataset.
@@
In this work, following [dp_spcom, dp_preprint, tit-preprint], we define and explicitly compute the worst-case error, over all datasets, of general preprocessing (or bounding) strategies. This error metric is natural in settings with arbitrarily correlated data, where a user potentially ascribes his/her error tolerance to the worst dataset that the statistic is computed on. Furthermore, this error metric is distribution-independent and, as we show, is computable under data heterogeneity too. We then explicitly identify the bounding strategy that results in the smallest worst-case error. Interestingly, we also observe from experimental studies that for scalar samples, our clipping strategy also gives rise to much smaller errors on average compared to that in [amin], for selected dataset sizes, when the samples are drawn i.i.d. according to common distributions.
@@ II-B Problem Formulation
Let $L$ be the number of users present in the dataset. For every user $\ell\in[L]$, let the number of records contributed by the user be denoted by $m_{\ell}$, and set $m^{\star}:=\max_{\ell\in[L]}m_{\ell}$ and $m_{\star}:=\min_{\ell\in[L]}m_{\ell}$. We assume that $L$ and the collection $\{m_{\ell}:\ell\in[L]\}$ are known. Now, for a given user $\ell\in[L]$, let $\left\{\mathbf{x}_{j}^{(\ell)}:j\in[m_{\ell}]\right\}$ denote the collection of (potentially arbitrary) bounded samples contributed by the user, where each $\mathbf{x}_{j}^{(\ell)}\in\mathbb{R}^{d}$, for some dimension $d\geq 1$. We assume, as is common for most applications [dpsgd, dp_spcom], that $\mathbf{x}_{j}^{(\ell)}\in\Delta_{U}$, for all $\ell\in[L]$, $j\in[m_{\ell}]$, where $U>0$ is known. Call the dataset as $\mathcal{D}=\left\{\left(\ell,\mathbf{x}_{j}^{(\ell)}\right):\ell\in[L],j\in[m_{\ell}]\right\}$.
We are interested in releasing the sample mean
$f=f(\mathcal{D}):=\frac{1}{\sum_{\ell^{\prime}=1}^{L}m_{\ell^{\prime}}}\cdot\sum_{\ell=1}^{L}\sum_{j=1}^{m_{\ell}}\mathbf{x}_{j}^{(\ell)}.$
@@ II-C Differential Privacy
[…]
@@
We use the terms “sensitivity” and “user-level sensitivity” interchangeably. The next result is well-known [dwork06, Prop. 1].
@@ II-D The Worst-Case Error Metric
All through, in this paper, we shall work with user-level $\varepsilon$-DP mechanisms that add a suitable amount of Laplace noise that is tailored to the sensitivity of the function used as an estimator of the sample mean $f$. Consider a mechanism $M$ for the user-level $\varepsilon$-DP release of the statistic $f$. The canonical structure of $M$ (see [staircase, Footnote 1], [opt]) is: $M(\mathcal{D})=\overline{f}(\mathcal{D})+{\mathbf{Z}},$ for some estimator $\overline{f}$ of $f$, with user-level sensitivity $\Delta_{\overline{f}}$, with $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$, with $Z_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\text{Lap}\left(\Delta_{\overline{f}}/\varepsilon\right)$, $i\in[d]$.
@@
$E_{M}=E_{\overline{f}}^{(d)}:=\max_{\mathcal{D}\in\mathsf{D}}\big\lVert f(\mathcal{D})-\overline{f}(\mathcal{D})\big\rVert_{1}+\mathbb{E}[\lVert{\mathbf{Z}}\rVert_{1}].$ (1)
Clearly, the expression above is an upper bound on the $\ell_{1}$-error $E^{(1)}_{M}:=\max_{\mathcal{D}\in\mathsf{D}}\mathbb{E}[|f(\mathcal{D})-M(\mathcal{D})|]$, via the triangle inequality. When the dimension $d$ is clear from the context, we simply denote $E_{\overline{f}}^{(d)}$ as $E_{\overline{f}}$. The distribution-independent expression in (1) conveniently captures the errors due to bias (the first term) and due to noise addition for privacy (the second term); a similar such error measure that separates the bias and noise errors was employed in [amin].
III Worst-Case Errors of Bounding/Clipping Strategies
@@ III-A On Clipping Strategies
We work with estimators $\overline{f}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ of $f$ obtained by bounding user contributions as follows: for each $\ell\in[L]$ and $j\in[m_{\ell}]$, we let $\overline{\mathbf{x}}_{j}^{(\ell)}:={\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}}\left(\mathbf{x}_{j}^{(\ell)}\right)$, for reals $0\leq a_{j}^{(\ell)}\leq b_{j}^{(\ell)}\leq U$. In words, $\overline{\mathbf{x}}_{j}^{(\ell)}$ is a projection of $\mathbf{x}_{j}^{(\ell)}$ onto the set $\mathsf{A}_{a_{j}^{(\ell)},b_{j}^{(\ell)}}$ (see Figure 1), which, intuitively, reduces the range of values that $\mathbf{x}_{j}^{(\ell)}$ can take, and hence its sensitivity too. We then set
@@
Note that the class $\mathsf{B}$ of estimators $\overline{f}$ as above captures those estimators obtained by dropping selected samples $\mathbf{x}_{j}^{(\ell)}$ (by setting $a_{j}^{(\ell)}=b_{j}^{(\ell)}=0$ for those samples) and those obtained by projecting samples onto an $\ell_{1}$-bounded subset of $\Delta_{U}$, in addition to strategies that perform a combination of dropping and projection. This class of estimators hence includes several common estimators of the sample mean used in works such as [userlevel, dp_preprint, tit-preprint].
@@ III-B On Worst-Case Errors of Clipping Strategies
Consider the quantity $E_{\overline{f}}^{(d)}$, for $\overline{f}\in\mathsf{B}$. The following proposition then holds:
@@
$E_{\overline{f}}^{(1)}$ $=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}$
and for any $d\geq 2$,
$E_{\overline{f}}^{(d)}$ $=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right)\ +$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{2d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
The proof of Proposition III.1 proceeds with help from a few lemmas. For $E_{\overline{f}}$ as in (1), we define $\beta^{(d)}(\overline{f})=\beta(\overline{f})$ to be the bias, i.e., $\beta(\overline{f}):=\max_{\mathcal{D}\in\mathsf{D}}\big\lVert f(\mathcal{D})-\overline{f}(\mathcal{D})\big\rVert$ and $\eta^{(d)}(\overline{f})=\eta(\overline{f})$ to be the error due to noise addition, i.e., $\eta(\overline{f}):=\mathbb{E}[\lVert\mathbf{\mathbf{Z}}\rVert_{1}]$. First, we aim to characterize $\beta(\overline{f})$. To this end, we first state a necessary condition for a vector $\mathbf{y}$ to be an $\ell_{1}$-projection of a vector $\mathbf{a}\in\Delta_{U}$ onto $\Delta_{\alpha}$, for $\alpha\leq U$.
@@
In the appendix, we explicitly characterize the vectors $\mathbf{y}\in\Delta_{\alpha}$ that are the $\ell_{1}$-projections of $\mathbf{a}\in\Delta_{U}\setminus\Delta_{\alpha}$, for $\alpha\leq U$; indeed, we have that the condition stated in Lemma III.1 is both necessary and sufficient. The next lemma exactly characterizes the bias $\beta^{(d)}(\overline{f})$, for any $d\geq 1$.
@@
We have that for any $d\geq 1$,
$\beta^{(d)}(\overline{f})=\frac{1}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}\cdot\left(\sum_{\ell\leq L}\sum_{j\leq m_{\ell}}\max\left\{a_{j}^{(\ell)},U-b_{j}^{(\ell)}\right\}\right).$
@@
We have that
$\eta^{(1)}(\overline{f})=\frac{d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}\left(b_{j}^{(\ell)}-a_{j}^{(\ell)}\right)}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}$
and for any $d\geq 2$,
$\eta^{(d)}(\overline{f})=\frac{2d\cdot\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\varepsilon\cdot\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}}.$
@@
Consider first the case when $d=1$. Via arguments entirely analogous to that in the proof of Lemma III.2, the user-level sensitivity of $\overline{f}$ is
@@
since in the worst-case, all samples of a user $\ell$ are changed each from $(a_{j}^{(\ell)},0,\ldots,0)$ to $(b_{j}^{(\ell)},0,\ldots,0)$. Next, for any $d\geq 2$, we have via similar arguments that
$\Delta_{\overline{f}}=\frac{2\max_{\ell\leq L}\sum_{j\leq m_{\ell}}b_{j}^{(\ell)}}{\sum_{\ell^{\prime}\leq L}m_{\ell^{\prime}}},$
since in the worst-case, all samples of a user $\ell$ are changed each from $(b_{j}^{(\ell)},0,0,\ldots,0)$ to $(0,b_{j}^{(\ell)},0,\ldots,0)$. Using $\mathbb{E}[\lVert\mathbf{Z}\rVert_{1}]=d\mathbb{E}[|Z_{1}|]=d\Delta_{\overline{f}}/\varepsilon$ gives us the required result. ∎
@@
We have that $f^{\star}=\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizes $E_{\overline{f}}$, where $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{2m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{Um_{\ell}+T_{\varepsilon}}{2m_{\ell}},U\right\},$ if $d=1$, and $a_{j}^{(\ell)}=\max\left\{\frac{Um_{\ell}-T_{\varepsilon}}{m_{\ell}},0\right\}\text{ and }b_{j}^{(\ell)}=\min\left\{\frac{T_{\varepsilon}}{m_{\ell}},U\right\}$, for any $d\geq 2$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.
Some remarks are in order. First, note that the optimal bounding strategy $f^{\star}$ clips sample values based only on the number of contributions of each user $\ell\in[L]$. Furthermore, note that the interval of projection is determined by $T_{\varepsilon}$, which is quite similar in structure to the optimal clipping threshold $T$ in [amin] for item-level DP, which is the (privately estimated) $\left\lceil\left(\frac{2}{\varepsilon}\right)\right\rceil^{\text{th}}$-largest sample value.
@@
For any $d\geq 1$, there exists an estimator $\overline{f}_{\left\{a_{j}^{(\ell)},b_{j}^{(\ell)}\right\}}$ minimizing $E_{\overline{f}}$, which obeys $a_{j}^{(\ell)}+b_{j}^{(\ell)}=U$, for all $\ell\in[L]$ and $j\in[m_{\ell}]$.